prove that 2^n > 1+n*√2^(n-1) for all n>2
1
Expert's answer
2019-03-07T10:33:44-0500


Prove that 2n > 1+n*√2n-1 for all n>2

Solution.

2n-1 > n*√2n-1 square both sides of inequality

(2n-1)2 > n2*2n-1

22n-2*2n+1 > n2*2n-1 move n2*2n-1 to the left side, than take 2n out of brackets

2n *( 2n - 2 – n2/2) +1 > 0

Obviously 2n > 0

So now we should just prove that ( 2n - 2 - n2/2) >= 0

2n - 2 >= n2/2

2(2n - 2) >= n2

As function graphs show

(where red graph – n2; blue graph - 2(2n - 2) and green one – n=2)

When n > 2 blue graph is higher than red one. That means that 2(2n - 2) >= n2.

And that proves that 2n > 1+n*√2n-1 for all n>2